4
For the reaction: Ca+H_2O rightarrow Ca(OH)_2 +H_2
Answer #1

7) Ca + 2H2O ----> Ca(OH)2 + H2

0.1 g Ca = 0.1g /40 g/mol = 0.0025 mole

So [OH-] = 2* 0.0025 moles / 0.500 L = 0.01 M

8) 0.1 M ZnCl2 contains 2*0.1 M (0.2 M)of Cl- ions

0.2 M = 0.2 mole / 1 L so in 1.75 L it will have 0.375 moles of Cl- ion